# FIAN/TD/12-2024

### v2

@@ Conclusion
In this note, we constructed an unfolded formulation of $4d$ pure Yang–Mills theory making use of the unfolding method proposed in [18] and improved here by preliminary derivation of general relations for unfolded functions.
A natural generalization of our result would be the inclusion of charged matter. It is straightforward and should not cause any problems, as shown by the example of scalar electrodynamics [18].
Another possible direction of further research is to include supersymmetry, as well as to manifest conformal symmetry. To this end one needs to introduce the corresponding gauge 1-forms of (super)conformal gravity in addition to the Poincaré connection and to deform appropriately the unfolded equations. In its turn, this requires further non-trivial modifications of the unfolding method of [18]. In particular, it would be interesting to unfold $\mathcal{N}=4$ super-Yang–Mills theory in order to apply unfolding tools to the problems of $AdS/CFT$ and amplitudes.
Quantization of the unfolded Yang–Mills system can be performed along the lines of [12] but with necessary modifications in order to include ghosts. The first step is to build an off-shell extension of the unfolded system constructed here, which is equivalent to coupling it to external currents [30, 31].
Finally, there are two vague but potentially promising topics: the study of integrability and the derivation of twistors. Both of them seem to be related to the investigation of various unfolding maps defined by the unfolded Yang–Mills equations.

### v3

@@ Conclusion
In this note, we constructed an unfolded formulation of $4d$ pure Yang–Mills theory making use of the unfolding method proposed in [21] and improved here by the preliminary derivation of general relations for unfolded functions.
A natural generalization of our result would be the inclusion of charged matter. It is straightforward and should not present any difficulties, as shown by the example of scalar electrodynamics [21].
Another possible direction of further research is to include supersymmetry, as well as to manifest conformal symmetry. To this end one needs to introduce the corresponding gauge 1-forms of (super)conformal gravity in addition to the Poincaré connection and to deform appropriately the unfolded equations. In its turn, this requires further non-trivial modifications of the unfolding method of [21]. In particular, it would be interesting to unfold $\mathcal{N}=4$ super-Yang–Mills theory in order to apply unfolding tools to the problems of $AdS/CFT$ and amplitudes. Another interesting problem is to unfold Einstein equations.
Quantization of the unfolded Yang–Mills system can be performed along the lines of [12] but with necessary modifications in order to include ghosts. The first step is to build an off-shell extension of the unfolded system constructed here, which is equivalent to coupling it to external currents [36, 37]. After that, this off-shell system can be elevated to an unfolded Schwinger–Dyson system for correlation functions of quantum unfolded fields.
Finally, there are two vague but potentially promising topics: the study of integrability of classical e.o.m. and the derivation of twistors. Both of them seem to be related to the investigation of various unfolding maps defined by the unfolded Yang–Mills equations.